OpenStudy (anonymous):

Express 0.40 x 0.50 as a fraction: 1/4 Not Here 1/3 1/5 1/6

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

thx superstar you been a great help

OpenStudy (anonymous):

thats what im here for

OpenStudy (anonymous):

=0.20 = 20/100 = 1/5

OpenStudy (anonymous):

if u post your next question on this post it will tell me when u post

OpenStudy (anonymous):

am done for know thx i'll bee look for you when i need it lol..

OpenStudy (anonymous):

im just that great right

OpenStudy (anonymous):

lol i guess so